OpenStudy (anonymous):

1. The table below shows the values of y for each value of x. X Y -1 0 1 -1 2 1 2 -2 Part A: Does the table represent a function? Justify your answer by using the values shown in the table No this is not a function because for it to be a function a value of x should pinpoint a value of y. Part B: The function f(x) shown below represents the number of fieldgoals Frank made at different distances, x, in meters, from the post. F(x) = -x + 8 Calculate and interpret the meaning for f(5). Part C: Write an ordered pair to represent the input and output of the function in Part

OpenStudy (anonymous):

@loser66

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

thats creepy

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

your pic. xD but i was kiding

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

I like my picture

OpenStudy (anonymous):

me too it be a kitty my fav. animal

OpenStudy (anonymous):

lol no its not i think its a picture of you...

OpenStudy (anonymous):

@hartnn please help

OpenStudy (anonymous):

@gorv someone help!

OpenStudy (texaschic101):

It is not a function. A function cannot have 2 or more x values that are the same. They can have 2 or more y values that are the same, just not the x ones. f(x) = -x + 8 f(5) = -5 + 8 f(5) = 3 In a function, the input is x and the output is y. f(x) is the same as y so we have : y = -x + 8 --- sub in any number you want for x and find y...lets try x = 4 y = -4 + 8 y = 4 One ordered pair is : (4,4)

OpenStudy (anonymous):

Ok thank you any idea for the rest?

OpenStudy (anonymous):

nvm i got the rest i just needed that thank you!!!